The average score of five students in a Mathematics quiz is 140. all of them scored above 138. what is the highest possible score among the students?
Answer by ikleyn(52905)

.

The maximum for one single student of five is achieved when four other students have the lowest possible grade of 139

under given conditions. 


Then this extraordinary student will have his (or her) outstanding maximum score of 140 + 4 = 144.



CHECK.  4*139 + 144 = 700 = 140*5.    ! Correct !

Solved.